OpenStudy (anonymous):

find the solution to this system of equations. 3x+7y=46 -3x-8y=-50

OpenStudy (anonymous):

we can use the elimination method 3x+7y=46 -3x-8y=-50 we are adding the equation to get rid of a variable, in this case x so -y=-4 divide by -1 y=4 then substitute y into either equation

OpenStudy (anonymous):

3x+7y=46 3x+7(4)=46 3x+28=46 3x=18 x=6

OpenStudy (anonymous):

thank you :)